is the following system classified as dependent, independent or inconsistent
-4x-2y=-4, 3x-y=13

Answer by stanbon(75887)

is the following system classified as dependent, independent or inconsistent
-4x-2y= -4
3x - y= 13
------------------
Modify:
-4x - 2y = -4
6x - 2y = 26
------------------
Subtract:
10x = 30
x = 3
-----
If x = 3, solve for "y":
3x - y = 13
9 - y = 13
y = -4
-----
The solution is (3,-4)
A single solution indicated the system is independent.
-------------------
No solution means inconsistent.
Infinitely many solutions means dependent.
================================
Cheers,
Stan H.